OpenStudy (anonymous):

Two numbers differ by 10 and have a product of 704. Which of the following equations could be used to find the two numbers?

OpenStudy (anonymous):

A. x(10-x)=704 B. x(10x)=704 C. x(x-10)=704 D. x+(x-10)=704

OpenStudy (nurali):

the two numbers are x and y. If they differ by 10, then x - y = 10, so y = x - 10. Now xy = 704, so x(x-10) = 704.

OpenStudy (compassionate):

(x - 10)x = 704 would be one.

OpenStudy (anonymous):

Okay, that's what I got the first time. Thanks for the clarification!